# How do Laws Produce the Future?

### v1

@@ Abstract
The view that the laws of nature produce later states of the universe from earlier ones (prominently defended by Maudlin) faces difficult questions as to how the laws produce the future and whether that production is compatible with special relativity. This article grapples with those questions, arguing that the concerns can be overcome through a close analysis of the laws of classical mechanics and electromagnetism. The view that laws produce the future seems to require that the laws of nature take a certain form, fitting what Adlam has called “the time evolution paradigm.” Making that paradigm precise, we might demand that there be temporally local dynamical laws that take properties of the present and the arbitrarily-short past as input, returning as output changes in such properties into the arbitrarily-short future. In classical mechanics, Newton’s second law can be fit into this form if we follow a proposal from Easwaran and understand the acceleration that appears in the law to capture how velocity (taken to be a property of the present and the arbitrarily-short past) changes into the arbitrarily-short future. The dynamical laws of electromagnetism can be fit into this form as well, though because electromagnetism is a special relativistic theory we might require that the laws meet a higher standard: linking past light-cone to future light-cone. With some work, the laws governing the evolution of the vector and scalar potentials, as well as the evolution of charged matter, can be put in a form that meets this higher standard.
@@ 1 Introduction
The dynamic production account of physical laws can be contrasted with two alternatives. [^1] There are other options. Adlam, (2022a), Chen & Goldstein, (2022), and Chen, (2024) discuss the universals account (Platonic reductionism), the powers account (Aristotelian reductionism), and the modal structure account. The first is the Mill-Ramsey-Lewis best systems account, according to which the laws do not truly govern what happens in nature, but instead merely offer concise and informative descriptions of all the things that happen throughout the entire history of the universe (including its full past and future). [^2] For an introduction to the best systems account, see Loewer, (1996); Lange, (2008). Put poetically, the laws are patterns in the mosaic that is our universe. The second is a recent alternative put forward by Chen & Goldstein, (2022) and Adlam, (2022a) according to which the laws govern without dynamic production by placing constraints on possible histories for the universe (constraints that may or may not look like laws of time evolution).
There is much that could be said about the advantages and disadvantages of these three competing accounts. Here, my focus will be on defending dynamic production from two serious challenges facing the account. First, one might wonder how the laws produce the future. Second, one might be worried as to whether that production is compatible with special relativity.
@@
Temporally local dynamical laws are not hard to find. In classical mechanics, Newton’s second law can be fit into the above form if we follow a proposal from Easwaran, (2014) and understand the acceleration that appears in the law to capture how velocity (taken to be a property of the present and the arbitrarily-short past) changes into the arbitrarily-short future. We will see that classical electromagnetism can also be formulated as a theory with temporally local dynamical laws, though there exist alternative formulations that are certainly not temporally local (the Wheeler-Feynman and retarded action-at-a-distance approaches).
@@
Let us now turn to the second challenge. Dorst, (forthcoming) has criticized the dynamic production account for being in tension with the special theory of relativity. One objection is that special relativity pushes us to a block universe theory where the past and future are just as real as the present and there is no work for the laws to do in producing the future. The future is already there. This objection will be addressed briefly in section 2. Dorst’s other objection is that we cannot say that dynamic production is occurring from one time slice to the next without privileging a particular way of carving the universe into simultaneity slices. Maudlin is willing to introduce such a preferred foliation because he believes that in quantum physics a preferred foliation is the most natural way to account for the experiments that illustrate Bell’s theorem. Maudlin may be right about how the quantum physics will shake out, but I still think that it is valuable to understand how relativistic dynamic production might work without a preferred foliation for two reasons. First, there are important classical theories that do not need a preferred foliation. Second, quantum theories that include parallel universes may not need such a foliation.
In a relativistic theory like classical electromagnetism, we can go beyond merely requiring temporally local dynamical laws that connect past and future and require that the dynamical laws be spatiotemporally local, connecting the past and future light-cones at every space-time point. This higher standard can be called “the relativistic time evolution paradigm.” With some work, we will see that the laws of electromagnetism can be put in this form. There are three types of laws that must be analyzed. First, there are the equations describing how charged matter produces the electromagnetic field. In the Lorenz gauge, these are the relativistic wave equations for the vector and scalar potentials. Such wave equations are standardly presented alongside causality theorems proving that what happens at a given point is fully determined by what happens on any slice of its past light-cone. The wave equations can be manipulated to make this causal structure manifest so that they describe how properties of the arbitrarily-small past light-cone change along the future light-cone. Second, there is the Lorentz force law giving the electromagnetic force on charged matter. This can be rewritten so that the force on a body at a location depends only upon properties of the arbitrarily-small past light-cone. Third, we need an equation describing how charged matter responds to forces. For a point charge, this would be the relativistic extension of Newton’s second law, where the force on the charge can be set equal to the rate at which the charge’s relativistic momentum (interpreted as a property of its arbitrarily-short past) changes into its arbitrarily-short future.
@@
The article proceeds as follows: The next section presents a precise statement of the time evolution paradigm that theories must fit for the dynamic production account of laws to be tenable. Section 3 explores multiple ways of fitting Newton’s classical theory of gravity into the time evolution paradigm, highlighting some challenges and choice points that will be relevant to the subsequent discussion of electromagnetism. Section 4 gives a simple way to cast electromagnetism into the time evolution paradigm. Section 5 puts forward the relativistic time evolution paradigm. Section 6 presents a way of fitting electromagnetism into the relativistic time evolution paradigm, showing that it is possible and leaving open whether there might be a better way of doing so. Section 7 looks ahead to more advanced physics and considers the prospects for interpreting quantum field theory and general relativity as theories of dynamic production.
@@ 2 The Time Evolution Paradigm
Physics has indeed come a long way since Newton, but one can debate whether that progress has occurred within—or gone beyond—“the Newtonian time-evolution picture.”
Adlam is not alone in presenting the time evolution paradigm as behind the times. The 2023 Foundational Questions Institute (FQxI) essay contest prompt began with this sentence:
@@
The prompt later asks: “Could it have been otherwise? Should it be otherwise now?” The implication is that this is an outdated mold into which physical theories need no longer be cast. I take the conservative view that this mold should be retained and that more effort should be put into fitting existing and future theories into the mold. It has served physics well and should not be abandoned lightly.
In Newton’s physics, the laws are deterministic. [^4] Although the laws of Newtonian physics are generally presented as deterministic, there are in fact some subtle and difficult questions as to whether they are truly deterministic that we will not explore here (Earman,, 2004; Wilson,, 2009; van Strien,, 2021). This is sometimes explained by saying that the laws allow only a single future given just the state at a particular moment, but here we must be careful. How does one specify the physical state at a moment? Albert, (2000, pg. 9–10) gives two requirements for such a specification. First, the instantaneous state should describe features of that moment alone. Second, a full set of instantaneous states across all times should fully specify everything that occurs. You might initially think that the masses, positions, and velocities of bodies should all be included in the instantaneous state, but Albert uses his first requirement to disqualify velocity. Velocity is the rate at which position changes and the velocity at a moment can only be determined by considering (arbitrarily small) temporal neighborhoods around the moment in question. Strictly speaking, it is not a property of that moment alone. Excluding velocity does not lead to a problem with the second requirement because the velocities at any time can be determined once the positions are specified at all times. If we agree with Albert and say that velocities are not included in instantaneous states, then the laws of Newton’s physics will certainly fail to determine a unique future given the state at a moment. But, the laws will give a unique future given the state of the world over an arbitrarily short time interval. Let us thus follow Albert, (2000, pg. 11) (and Arntzenius,, 2000, pg. 195) and say that a theory is deterministic if and only if specifying the state of the world over an arbitrarily small time interval to the past of a given moment (assuming the laws to be obeyed during this interval) uniquely determines a single future that is allowed by the laws (a single future that is physically possible). [^5] Builes & Impagnatiello, (forthcoming) call this kind of determinism “Near Markovian Determination.”
@@
as in Lange, (2005, sec. 2). This velocity is depicted in figure 1. Modifying a proposal from Easwaran, (2014), [^6] Easwaran, (2014) uses clever, but more complicated, definitions of the past and future time derivatives. These definitions avoid referencing the present moment by either considering two arbitrarily close past times (for the past time derivative) or two arbitrarily close future times (for the future derivative). Easwaran calls these open-ended derivatives (because the endpoint at the present is left out) and contrasts them with the closed-ended derivatives defined above. Easwaran’s choice to use open-ended derivatives allows him to classify past velocity as entirely about the past and future acceleration as entirely about the future. One reason behind the choice to use closed-ended derivatives here will be given later (in footnote 31) when we see that, in electromagnetism, even if we were to use open-ended derivatives the laws would not give pure properties of the future light-cone as output. the acceleration that appears in the laws can be defined as the future acceleration $\vec{a}^{\,pf}$: a future derivative of the past velocity,
@@
Figure 2: The future acceleration $\vec{a}^{\,pf}$ (2) of a body is determined by comparing (a) the body’s past velocity at a time $\epsilon$ in the future (which is determined by looking $\delta$ into the past of that moment, as shown on the left), and (b) the body’s past velocity at the moment in question (which is determined by looking $\delta$ into the past of that moment, as shown on the right). As is evident from the image on the right, a body’s future acceleration $\vec{a}^{\,pf}$ depends on its past and is not a pure property of its present and future.
@@
A past neighborhood property of an object at $t$ is not determined by the present properties of the object at $t$ alone, but is determined once the present properties of the object are specified within any arbitrarily small interval $[t-\Delta,t]$. [^8] Because we need at least some of the past to determine the velocity (or, using other words, to serve as grounds for the velocity), but we do not need any particular past moment, there is a puzzle as to exactly what are the facts about position in virtue of which the velocity at a moment is what it is (Builes & Teitel,, 2019).
@@
Note that, by this definition, future neighborhood properties also count as past-to-future neighborhood properties because they are determined by the present properties over an arbitrarily small interval to the future. However, some past-to-future properties—like the acceleration $\vec{a}^{\,pf}$—are not future neighborhood properties (see figure 2). If we require dynamic laws to give past-to-future neighborhood properties as output, we can allow for either first-order dynamics (as in the Schrödinger equation) where the output is a future neighborhood property (and thus a past-to-future neighborhood property) or second-order dynamics (as in Newton’s second law) where the output is a past-to-future neighborhood property (that is not a future neighborhood property). [^9] One might also wonder about third and higher order dynamics. Easwaran, (2014) gives an argument against the possibility of such dynamics (cf. Swanson,, 2022).
To avoid action at a temporal distance, Easwaran, (2014, pg. 857) conjectures that “The fundamental causal [dynamical] laws must use present properties and past neighbourhood properties to determine future neighbourhood properties.” This idea has not been widely adopted, but it is an attractive proposal that fits well with a dynamic production account of laws. Let us take a modified version of the idea onboard here and say that a law counts as a temporally local dynamical law if it gives past-to-future neighborhood properties as output and takes as input only present properties or past neighborhood properties. Dynamical laws that are not temporally local might take as input properties at earlier or later times, as in the retarded action-at-a-distance and Wheeler-Feynman versions of electromagnetism that will be discussed in section 4. Dynamical laws that are temporally local may either take as input only present properties (as in the first-order Schrödinger equation) or both present and past neighborhood properties (as in Newton’s second law).
@@
    Temporally Local Dynamical Laws: A subset of the laws (the dynamical laws) can be applied at any moment, $t$, giving past-to-future neighborhood properties as output and taking as input only present properties or past neighborhood properties. The dynamical laws may give either precise values for the past-to-future neighborhood properties or probability distributions over such values.
@@
This formulation allows for non-dynamical laws in addition to dynamical laws. As a candidate example of a non-dynamical law, Maudlin, (2007, pg. 13) gives Newton’s law of gravitation, which uses present positions and masses to specify the present force on a given body. Maudlin describes this law as an “adjunct principle” that allows the dynamical law $\vec{F}=m\vec{a}$ to predict future time evolution. Chen & Goldstein, (2022, pg. 60) consider the example of Gauss’s law, $\vec{\nabla}\cdot\vec{E}=4\pi\rho$, which relates the present divergence of the electric field to the present distribution of charge. The requirement in the second condition that other laws “express relations between present properties or past neighborhood properties” rules out certain putative non-dynamical laws, such as those that posit probability distributions over initial conditions (Chen & Goldstein,, 2022, sec. 3.3.3). [^10] One could still posit a probability distribution over initial conditions to explain thermodynamic asymmetries, as Albert, (2000) does with his Past Hypothesis and Statistical Postulate. Within the time evolution paradigm, such a probability distribution would have to be regarded as a hypothesis (or postulate), not a law of nature. The third condition, that the laws be either deterministic or stochastic, is meant to exclude laws that give incomplete stories about the future: neither determining the future uniquely nor giving definite probabilities for different futures. As an example of such an incomplete theory, consider a version of electromagnetism where the vector and scalar potentials are taken to be the proper way to represent the electromagnetic field and no gauge fixing condition is imposed. Then, multiple future evolutions of the potentials will be compatible with the past (Maudlin,, 2018, pg. 14).
The dynamic production account of laws appears to require the laws of nature to fit the time evolution paradigm presented above. [^11] Relaxing the requirement of temporal locality, one might wonder whether laws that take more of the past as input could plausibly be construed as dynamically producing the future (as in, e.g., a retarded action-at-a-distance formulation of electromagnetism). Laws like that are indeed compatible with the broad idea of dynamic production, but they do not look like the FLOTEs that Maudlin envisages and I will take them to be at odds with the type of dynamic production account of laws that Maudlin has defended (and that I am defending here). Let us begin with the first condition of the time evolution paradigm. The temporally local dynamical laws that appear in this condition can serve as what Maudlin, (2007) calls the fundamental laws of temporal evolution (FLOTEs). It is these laws that generate time evolution, producing future states from earlier ones. As Maudlin puts it,
@@
Chen & Goldstein, (2022, pg. 46 & 60) challenge the idea of dynamic production because the state of the universe at a moment does not contain features like momentum that are needed to evolve the state forward via the laws. However, we can take momentum to be a past neighborhood property—like the past velocity in (1)—and understand the relation of dynamic production to connect the present, and its past neighborhood, to the future. The way that the dynamical laws in the time evolution paradigm connect past to future also helps to address Loewer,’s (2012, pg. 133) remark that “there isn’t a ‘next’ state if time is continuous” to support the claim that one state of the universe produces the next via the laws. The laws give as output a set of past-to-future-neighborhood properties that specify how things change into the future, not the state at the next moment.
@@
The third condition of the time evolution paradigm requires the laws to be either deterministic or stochastic. This is a natural requirement for the dynamic production account. To dynamically produce the future, the laws must be sufficiently specific that they either yield a unique future or a probability distribution over possible futures. [^15] If the laws give probabilities, those probabilities should be understood as basic propensities that are not to be reduced to anything more fundamental, such as long-run frequencies. See Maudlin, (2007, pg. 17–20); Frigg & Hoefer, (2007); Loewer, (2012, pg. 118).

### v2

@@ Abstract
The view that the laws of nature produce later states of the universe from earlier ones (prominently defended by Maudlin) faces difficult questions as to how the laws produce the future and whether that production is compatible with special relativity. This article grapples with those questions, arguing that the concerns can be overcome through a close analysis of the laws of classical mechanics and electromagnetism. The view that laws produce the future seems to require that the laws of nature take a certain form, fitting what Adlam has called “the time evolution paradigm.” Making that paradigm precise, we might demand that there be temporally local dynamical laws that take properties of the present and the arbitrarily-short past as input, returning as output changes in such properties into the arbitrarily-short future. In classical mechanics, Newton’s second law can be fit into this form if we follow a proposal from Easwaran and understand the acceleration that appears in the law to capture how velocity (taken to be a property of the present and the arbitrarily-short past) changes into the arbitrarily-short future. The dynamical laws of electromagnetism can be fit into this form as well, though because electromagnetism is a special relativistic theory we might require that the laws meet a higher standard: linking past light-cone to future light-cone. With some work, the laws governing the evolution of the vector and scalar potentials in the Lorenz gauge, as well as the evolution of charged matter, can be put in a form that meets this higher standard.
@@ 1 Introduction
The dynamic production account of physical laws can be contrasted with two alternatives. [^1] There are other options. Adlam, (2022a), Chen & Goldstein, (2022), and Chen, (2024) discuss the universals account (Platonic reductionism), the powers account (Aristotelian reductionism), and the modal structure account. The first is the Humean Mill-Ramsey-Lewis best systems account, according to which the laws do not truly govern what happens in nature, but instead merely offer concise and informative descriptions of all the things that happen throughout the entire history of the universe (including its full past and future). [^2] See Loewer, (1996); Lange, (2008). Put poetically, the laws are patterns in the mosaic that is our universe. The second is a recent alternative, put forward by Chen & Goldstein, (2022), Adlam, (2022a), and Meacham, (2022, forthcoming), according to which the laws govern without dynamic production by placing constraints on possible histories for the universe (constraints that may or may not look like laws of time evolution).
There is much that could be said about the advantages and disadvantages of these three competing accounts. Here, my focus will be on developing the dynamic production and defending it from two serious challenges. First, one might wonder how the laws produce the future. Second, one might be worried as to whether that production is compatible with special relativity.
@@
Temporally local dynamical laws are not hard to find. In classical mechanics, Newton’s second law can be fit into the above form if we follow a proposal from Easwaran, (2014) and understand the acceleration that appears in the law to capture how velocity (taken to be a property of the present and the arbitrarily-short past) changes into the arbitrarily-short future. Classical electromagnetism can also be formulated as a theory with temporally local dynamical laws, though there exist alternative formulations that are certainly not temporally local (the Wheeler-Feynman and retarded action-at-a-distance approaches).
@@
Let us now turn to the second challenge. Dorst, (forthcoming) has criticized the dynamic production account for being in tension with the special theory of relativity. One objection is that special relativity pushes us to a block universe theory where the past and future are just as real as the present and there is no work for the laws to do in producing the future. The future is already there. This objection will only be addressed briefly in section 2. Dorst’s other objection is that we cannot say that dynamic production is occurring from one time slice to the next without privileging a particular way of carving the universe into simultaneity slices. Maudlin is willing to introduce such a preferred foliation because he believes that, in quantum physics, a preferred foliation is the most natural way to account for the experiments that illustrate Bell’s theorem. Maudlin may be right about how the quantum physics will shake out, but I still think that it is valuable to understand how relativistic dynamic production might work without a preferred foliation for two reasons. First, there are important classical theories that do not need a preferred foliation. Second, quantum theories that include parallel universes may not need such a foliation.
In a relativistic theory like classical electromagnetism, we can go beyond merely requiring temporally local dynamical laws that connect past and future and require that the dynamical laws be spatiotemporally local, connecting the past and future light-cones at every space-time point. This higher standard, which can be called “the relativistic time evolution paradigm,” supports an understanding of dynamic production that does not require a preferred foliation.
@@
The article proceeds as follows: The next section presents a precise statement of the time evolution paradigm that theories must fit for the dynamic production account of laws to be tenable. Section 3 explores multiple ways of fitting Newton’s classical theory of gravity into the time evolution paradigm, highlighting some challenges and choice points that will be relevant to the subsequent discussion of electromagnetism. Section 4 gives a simple way to cast electromagnetism into the time evolution paradigm. Section 5 puts forward the relativistic time evolution paradigm. Section 6 explores (in depth) how the laws of electromagnetism might be fit into the relativistic time evolution paradigm, showing that it is possible and leaving open whether there might be a better way of doing so. Section 7 looks ahead to more advanced physics and considers the prospects for interpreting quantum field theory and general relativity as theories of dynamic production.
@@ 2 The Time Evolution Paradigm
Adlam is not alone in presenting the “Newtonian time-evolution picture” as behind the times. The 2023 Foundational Questions Institute (FQxI) essay contest prompt began with this sentence:
@@
The prompt later asks: “Could it have been otherwise? Should it be otherwise now?” The implication is that this is an outdated mold into which physical theories need no longer be cast. I take the conservative view that this mold should be retained, and that more effort should be put into fitting existing and future theories into the mold. It has served physics well and should not be abandoned lightly.
In Newton’s physics, the laws are deterministic. [^4] At least, the laws of Newtonian physics are generally presented as deterministic. The question as to whether the laws truly are deterministic will be raised near the end of this section. This is sometimes explained by saying that the laws allow only a single future given just the state at a particular moment, but here we must be careful. How does one specify the physical state at a moment? Albert, (2000, pg. 9–10) gives two requirements for such a specification. First, the instantaneous state should describe features of that moment alone. Second, a full set of instantaneous states across all times should fully specify everything that occurs. You might initially think that the masses, positions, and velocities of bodies should all be included in the instantaneous state, but Albert uses his first requirement to disqualify velocity. Velocity is the rate at which position changes and the velocity at a moment can only be determined by considering (arbitrarily small) temporal neighborhoods around the moment in question. Strictly speaking, it is not a property of that moment alone. Excluding velocity does not lead to a problem with the second requirement because the velocities at any time can be determined once the positions are specified at all times. If we agree with Albert and say that velocities are not included in instantaneous states, then the laws of Newton’s physics will certainly fail to determine a unique future given only the state at a moment. But, the laws will give a unique future given the state of the world over an arbitrarily short time interval. Let us thus follow Albert, (2000, pg. 11) (and Arntzenius,, 2000, pg. 195) and say that a theory is deterministic if and only if specifying the state of the world over an arbitrarily small time interval to the past of a given moment (assuming the laws to be obeyed during this interval) uniquely determines a single future that is allowed by the laws (a single future that is physically possible). [^5] Builes & Impagnatiello, (forthcoming) call this kind of determinism “Near Markovian Determination.”
@@
as in Lange, (2005, sec. 2). This velocity is depicted in figure 1. The future velocity, by contrast, would be defined by considering arbitrarily small time intervals to the future. In ordinary circumstances, the past and future velocities will agree. However, they could in principle come apart if there is a kink in the trajectory of a body through space and time. Whether past and future velocities are permitted to come apart will depend on the laws of the theory at hand.
Modifying a proposal from Easwaran, (2014), [^6] Easwaran, (2014) uses clever, but more complicated, definitions of the past and future time derivatives. These definitions avoid referencing the present moment by either considering two arbitrarily close past times (for the past time derivative) or two arbitrarily close future times (for the future derivative). Easwaran calls these open-ended derivatives (because the endpoint at the present is left out) and contrasts them with the closed-ended derivatives defined above. Easwaran’s choice to use open-ended derivatives allows him to classify past velocity as entirely about the past and future acceleration as entirely about the future. One reason behind the choice to use closed-ended derivatives here will be given later (in footnote 33) when we see that, in electromagnetism, even if we were to use open-ended derivatives the laws would not give pure properties of the future light-cone as output. the acceleration that appears in the laws can be defined as the future acceleration $\vec{a}^{\,pf}$: a future derivative of the past velocity,
@@
Figure 2: The future acceleration $\vec{a}^{\,pf}$ (2) of a body is determined by comparing (a) the body’s past velocity at a time $\epsilon$ in the future (which is determined by looking $\delta$ into the past of that moment, as shown on the left), and (b) the body’s past velocity at the moment in question (which is determined by looking $\delta$ into the past of that moment, as shown on the right). As is evident from the image on the right, a body’s future acceleration $\vec{a}^{\,pf}$ depends on its past and is not purely a property of its present and future.
@@
A past neighborhood property of an object at $t$ is not determined by the present properties of the object at $t$ alone, but is determined once the present properties of the object are specified within any arbitrarily small interval $[t-\Delta,t]$. [^8] Because we need at least some of the past to determine the velocity, but we do not need any particular past moment, there is a puzzle as to exactly what are the facts about position in virtue of which the velocity at a moment is what it is (Builes & Teitel,, 2019).
@@
Note that, by this definition, future neighborhood properties also count as past-to-future neighborhood properties because they are determined by the present properties over an arbitrarily small interval to the future. However, some past-to-future properties—like the acceleration $\vec{a}^{\,pf}$—are not future neighborhood properties (see figure 2). If we require dynamical laws to give past-to-future neighborhood properties as output, we can allow for either first-order dynamics (as in the Schrödinger equation) where the output is a future neighborhood property (and thus a past-to-future neighborhood property) or second-order dynamics (as in Newton’s second law) where the output is a past-to-future neighborhood property (that is not a future neighborhood property). [^9] One might also wonder about third and higher order dynamics. Easwaran, (2014) gives an argument against the possibility of such dynamics (cf. Swanson,, 2022).
To avoid action at a temporal distance, Easwaran, (2014, pg. 857) conjectures that “The fundamental causal [dynamical] laws must use present properties and past neighbourhood properties to determine future neighbourhood properties.” This idea has not been widely adopted, but it is an attractive proposal that fits well with a dynamic production account of laws. Let us take a modified version of the idea onboard here and say that a law counts as a temporally local dynamical law if it gives past-to-future neighborhood properties as output and takes as input only present properties or past neighborhood properties. Dynamical laws that are not temporally local might take as input properties at earlier or later times, as in the retarded action-at-a-distance and Wheeler-Feynman versions of electromagnetism. Dynamical laws that are temporally local may either take as input only present properties (as in the first-order Schrödinger equation) or both present and past neighborhood properties (as in Newton’s second law).
@@
    Temporally Local Dynamical Laws: A subset of the laws (the dynamical laws) can be applied at any moment $t$, take as input only present properties or past neighborhood properties, and give as output only past-to-future neighborhood properties. The dynamical laws may give either precise values for the past-to-future neighborhood properties or probability distributions over such values.
@@
This formulation allows for non-dynamical laws in addition to dynamical laws. As a candidate example of a non-dynamical law, Maudlin, (2007, pg. 13) gives Newton’s law of gravitation, which uses present positions and masses to specify the present force on a given body. Maudlin describes this law as an “adjunct principle” that allows the dynamical law $\vec{F}=m\vec{a}$ to predict future time evolution. Chen & Goldstein, (2022, pg. 60) consider the example of Gauss’s law, $\vec{\nabla}\cdot\vec{E}=4\pi\rho$, which can be regarded as a non-dynamical law relating the present divergence of the electric field to the present distribution of charge. The requirement in the second condition that non-dynamical laws “express relations between present properties or past neighborhood properties” rules out certain putative non-dynamical laws, such as those that posit probability distributions over initial conditions (Chen & Goldstein,, 2022, sec. 3.3.3). [^10] One could still posit a probability distribution over initial conditions to explain thermodynamic asymmetries, as Albert, (2000) does with his Past Hypothesis and Statistical Postulate. Within the time evolution paradigm, such a probability distribution would have to be regarded as a hypothesis (or postulate), not a law of nature. The third condition, that the laws be either deterministic or stochastic, is meant to exclude laws that give incomplete stories about the future: neither determining the future uniquely nor giving definite probabilities for different futures. As an example of such an incomplete theory, consider a version of electromagnetism where the vector and scalar potentials are taken to be the proper way to represent the electromagnetic field and no gauge fixing condition is imposed. Then, multiple future evolutions of the potentials will be compatible with the past (Maudlin,, 2018, pg. 14).
The dynamic production account of laws appears to require the laws of nature to fit the time evolution paradigm presented above. Let us begin with the first condition of the time evolution paradigm. The temporally local dynamical laws that appear in this condition can serve as what Maudlin, (2007) calls the fundamental laws of temporal evolution (FLOTEs). [^11] Relaxing the requirement of temporal locality, one might wonder whether laws that take more of the past as input could plausibly be construed as dynamically producing the future (as in, e.g., a retarded action-at-a-distance formulation of electromagnetism). Laws like that are indeed compatible with the broad idea of dynamic production, but they do not look like the FLOTEs that Maudlin envisages and I will take them to be at odds with the type of dynamic production account of laws that Maudlin has defended (and that I am defending here). It is these laws that generate time evolution, producing future states from earlier ones. As Maudlin puts it,
@@
Chen & Goldstein, (2022, pg. 46 & 60) challenge the idea of dynamic production because the state of the universe at a moment does not contain features like velocity and momentum that are needed to evolve the state forward via the laws. However, velocity and momentum can be formulated as past neighborhood properties and such properties may serve as inputs to temporally local dynamical laws connecting the present and its past neighborhood to the future. The way that the dynamical laws in the time evolution paradigm connect past to future also helps to address Loewer,’s (2012, pg. 133) remark that “there isn’t a ‘next’ state if time is continuous” to support the claim that one state of the universe produces the next via the laws. The laws give as output a set of past-to-future-neighborhood properties that specify how things change into the future, not the state at the next moment.
@@
The third condition of the time evolution paradigm requires the laws to be either deterministic or stochastic. This is a natural requirement for the dynamic production account and it is one that is imposed by Maudlin, (2007, pg. 14–20, 174–183). To dynamically produce the future, the laws must be sufficiently specific that they either yield a unique future or a probability distribution over possible futures. In a deterministic theory, the dynamical laws must together say enough about past-to-future neighborhood properties to specify a single future evolution—when applied at all times, alongside the non-dynamical laws. This is the simplest way that the laws might produce the future and a narrow version of the dynamic production account might require the laws to take this form. By contrast, stochastic laws allow for multiple futures and give probabilities for each. [^15] These probabilities should be understood as basic propensities that are not to be reduced to anything more fundamental, such as long-run frequencies. See Maudlin, (2007, pg. 17–20); Frigg & Hoefer, (2007); Loewer, (2012, pg. 118). Such laws can arguably also be regarded as producing the future, though they do so in a fundamentally random way where different futures have different chances of occurring. One can write down laws that are neither deterministic nor stochastic, deeming multiple futures to be compatible with a given past. It is hard to see how such laws could be understood as producing the future. Such laws do not specify, even in a probabilistic way, which future is to be produced (leaving open multiple possibilities). Here are three case where that might occur: First, as was mentioned previously, this kind of indeterminism arises if electromagnetism is formulated in terms of potentials without a gauge fixing condition. Maudlin, (2018) advocates adopting a gauge-fixing condition to avoid this indeterminism. Second, this kind of indeterminism may also occur in general relativity if one allows for the possibility of closed time-like curves (Earman et al.,, 2024). Maudlin, (2007, pg. 175, 189–191) forbids closed time-like curves as incompatible with his picture of dynamic production. Third, there have been worries raised about Newton’s laws of classical mechanics leaving open multiple futures and thus not being as deterministic as they appear (Earman,, 2004; Wilson,, 2009; van Strien,, 2021). A defender of dynamic production would seek sets of laws that avoid such indeterminism.
Unlike the dynamic production account, the best systems account does not require laws to fit the time evolution paradigm. Still, deterministic or stochastic temporally local dynamical laws have a good shot at being included in the best system because they get you a lot from a little—allowing you to predict the future (at least probabilistically) from an arbitrarily thin time slice. Callender, (2017, ch. 7-8) argues that time is a special dimension within the best systems account because the laws that emerge as part of the best system include laws of time evolution. As Callender, (2017, pg. 142) puts it, “Time is that direction on the manifold of events in which we can tell the strongest or most informative stories.” Although the laws of the best system might include dynamical laws, the role of these laws is different than in the dynamic production account. The best systems account starts with a history of things that happen and then finds laws as patterns in that history. By contrast, the dynamic production account views the history as generated by the laws. The laws are additional features of the universe beyond the things within the universe that are governed by the laws.
A defender of dynamic production may allow that it is metaphysically possible to have laws of nature that violate the time evolution paradigm and cannot be understood as dynamically producing the future from the past. Adopting this perspective, dynamic production is not part of what it takes for something to be a law. It is merely a feature of our laws. It is difficult to say what it takes in general for something to be a law, but you might say that at a minimum it must constrain what is physically possible. Here the dynamic production account can make peace with constraint accounts (Chen & Goldstein,, 2022; Adlam,, 2022a; Meacham,, 2022, forthcoming). Proponents of constraint accounts prefer the flexibility of their picture, arguing that the laws of our world may well turn out not to include dynamical laws (or to otherwise violate the time evolution paradigm) and that our philosophical theorizing should not preclude this possibility. [^16] In addition to Chen & Goldstein, (2022); Adlam, (2022a); Meacham, (2022, forthcoming), see also Wharton, (2015); Lam & Wüthrich, (2023). Proponents of dynamic production stick their necks out, betting that the final laws will fit the time evolution paradigm because our most successful extant theories are theories of time evolution (or at least can be formulated as theories of time evolution). This debate hinges primarily on difficult questions about the interpretation of general relativity and quantum field theory that will only be discussed briefly in section 7. Here we focus on theories that seem like they should fit the time evolution paradigm, to better understand how dynamic production works in these cases.